differntiate: y= xlog 10 x y= x-1/log 2 x

differntiate:


y= xlog10 x


y= x-1/log2


Yell 

Grade:11

2 Answers

Godfrey Classic Prince
12 years ago

Dear Shreya Chandwadkar,

In the first question use the UV formula ...

UV=UV/ + VU/

I am getting the answer to be ...

y=x*1 / x ln10 + log10 x

In the second question use the U/V formula...

U/V = VU/-UV/ / V2

I am getting the answer to be ..

y= [log2x - (x-1)*1/xln2] / (xlog2)2

Please simplfy the second answer ...for your practice dear !!

Hope this helped you...

If you loved the answer then please do not forget to approve it by clicking on "Yes" below..!!!

Abhishek Panhale
12 years ago

y= x logv10 x

dy/dx= x * logx/log10

        =1/log10[ x d(logx)/dx + logx d(x)/dx]

        =1/log10[ x/x + logx]

        =1/log10[ 1 + logx]

        =1/log10 + logv10 x

as simple as dat!!!!